# Black holes immersed in modified Chaplygin-like dark fluid and cloud of strings: shadows, quasinomal modes and greybody factors

### v1

@@ I Introduction
These observational breakthroughs demand a deeper theoretical understanding of black holes in environments that more realistically reflect their cosmological context. To this end, three principal theoretical tools have emerged as crucial diagnostics of black hole structure and dynamics. Black hole shadows characterize the optical appearance through critical photon orbits [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Recent explorations have expanded this framework to diverse environments [21, 22, 23, 24, 25], with contemporary research even investigating the intriguing possibility of using shadows as novel probes for particle dark matter [26]. QNMs reveal the stability and characteristic ringing of spacetime under external perturbations [27, 28, 29, 30, 31, 32, 33, 34, 35]. Greybody factors (GBFs) govern the transmission probability of Hawking radiation to distant observers, providing essential insights into semi-classical evaporation processes and quantum effects in curved spacetime [36, 37, 38]. Recent studies have revealed profound connections among these three diagnostics. In the eikonal limit, a well-established correspondence links QNM frequencies to the properties of the photon sphere and consequently to the shadow radius [39, 40, 41, 42, 43, 44, 45]. Furthermore, the interrelationship between shadows, QNMs, and GBFs has solidified into a unified paradigm, providing a unified paradigm to synergistically probe strong-field gravity and exotic environmental effects [46, 47, 48, 49, 50]. These interconnections provide a unified framework for understanding black hole physics across different observational channels.
@@
Further enriching this landscape, the cloud of strings (CoS) model describes how extended classical objects can modify black hole spacetimes. As one of the simplest extensions beyond point particles, a CoS alters both the metric and thermodynamic properties of black holes, offering insights into how fundamental string-theoretic structures might manifest in classical gravity [80]. Recently, studies of black holes immersed in a CoS have attracted growing interest [59, 24, 81].
@@
Figure 1: The metric function $f(r)$ for diverse parameter configurations. The first panel displays an MCDF-CoS reference case with $a=0.2$, $A=1$, $B=10^{-5}$, $\beta=0.8$, $Q=0.5$, alongside specific limits: MCDF ($a=0$), GCDF ($a=0$, $A=0$, $\beta\neq 1$), CDF ($a=0$, $A=0$, $\beta=1$), and Schwarzschild-de Sitter ($a=0$, $Q=0$). Subsequent panels vary individual parameters while maintaining others at reference values.
@@ II Black hole immersed in a modified Chaplygin-like dark fluid and a cloud of strings
When $a=0$, the black hole metric (1) describes a black hole immersed solely in an MCDF. In the limit $Q\rightarrow 0$, the MCDF metric reduces to
@@ IV.3 QNM results and discussion
Our systematic analysis of MCDF-CoS QNMs, computed via third-order WKB and Pöschl-Teller methods (Tables 14, Figs. 46), reveals strong correlations with the effective potential: higher barriers increase $\mathrm{Re}(\omega)$ (oscillation frequency), while broader barriers decrease $|\mathrm{Im}(\omega)|$ (damping rate).
@@
The overtone structure (Table 3, Fig. 6) shows higher overtones ($n>0$) exhibit larger $|\mathrm{Im}(\omega)|$ (faster decay), consistent with generic black hole behavior. For fundamental modes, $|\mathrm{Im}(\omega)|$ decreases and asymptotes with increasing $l$, while $\mathrm{Re}(\omega)$ grows approximately linearly.
@@
Throughout our analysis, the third-order WKB and Pöschl-Teller methods show excellent agreement for low-overtone fundamental modes, validating our computational approach.
@@
Table 3: Dependence of QNM frequencies on the angular quantum number $l$ and overtone number $n$ for the reference black hole configuration: $a=0.2$, $A=1$, $B=10^{-5}$, $\beta=0.8$, $Q=0.5$. $\Delta_{R}\%$ and $\Delta_{I}\%$ denote the relative errors for real and imaginary parts, calculated as $|\omega_{\mathrm{WKB}}-\omega_{\mathrm{P-T}}|/|\omega_{\mathrm{WKB}}|\times 100\%$.
@@
5 2 0.734196 - 0.307747 $\mathrm{i}$ 0.750694 - 0.305754 $\mathrm{i}$ 2.25 0.65
5 3 0.720096 - 0.433797 $\mathrm{i}$ 0.750694 - 0.428056 $\mathrm{i}$ 4.25 1.32
5 4 0.702849 - 0.561791 $\mathrm{i}$ 0.750694 - 0.550357 $\mathrm{i}$ 6.80 2.04
@@
6 2 0.872765 - 0.306945 $\mathrm{i}$ 0.886885 - 0.305480 $\mathrm{i}$ 1.62 0.48
6 3 0.860402 - 0.432006 $\mathrm{i}$ 0.886885 - 0.427672 $\mathrm{i}$ 3.08 1.00
6 4 0.845028 - 0.558747 $\mathrm{i}$ 0.886885 - 0.549865 $\mathrm{i}$ 4.95 1.59
6 5 0.827159 - 0.687061 $\mathrm{i}$ 0.886885 - 0.672057 $\mathrm{i}$ 7.22 2.18
@@
Figure 8: QNM-GBF correspondence for MCDF-CoS black holes with parameters $a=0.2$, $A=1$, $B=10^{-5}$, $\beta=0.8$, $Q=0.5$. From left to right, the multipole number l=1,10,100l=1,~10,~100, respectively.
@@ V.1 Eikonal limit connection
For massless scalar perturbations, the GBF can be expressed in terms of the fundamental QNM frequency $\omega_{0}$ through the relation [46]:
$\Gamma_{l}(\omega)\approx\frac{1}{1+\exp\left[2\pi\frac{\omega^{2}-\mathrm{Re}(\omega_{0})^{2}}{4\mathrm{Re}(\omega_{0})\mathrm{Im}(\omega_{0})}\right]},$ (45)
where $\omega_{0}$ denotes the complex QNM frequency computed via the WKB approximation.
Employing the eikonal QNM approximation (41) for fundanmental modes yields
@@
A notable observation is that while the QNM-GBF correspondence in Eq. (45) demonstrates excellent numerical agreement across the parameter space (Fig. 8), the shadow-GBF relation in Eq. (47) exhibits more pronounced discrepancies. This enhanced sensitivity likely stems from the exponential amplification of small deviations between the actual QNM real part and its geometric optics approximation ($\mathrm{Re}(\omega_{0})-\Omega_{\text{ps}}l$) within the GBF expression. Nevertheless, both formulations consistently underscore the fundamental role of photon sphere geometry in governing wave scattering dynamics.

### v2

@@ I Introduction
These observational breakthroughs demand a deeper theoretical understanding of black holes in environments that more realistically reflect their cosmological context. To this end, three principal theoretical tools have emerged as crucial diagnostics of black hole structure and dynamics. Black hole shadows characterize the optical appearance through critical photon orbits [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Recent explorations have expanded this framework to diverse environments [25, 26, 27, 28, 29], with contemporary research even investigating the intriguing possibility of using shadows as novel probes for particle dark matter [30]. QNMs reveal the stability and characteristic ringing of spacetime under external perturbations [31, 32, 33, 34, 35, 36, 37, 38, 39]. Greybody factors (GBFs) govern the transmission probability of Hawking radiation to distant observers, providing essential insights into semi-classical evaporation processes and quantum effects in curved spacetime [40, 41, 42].
Furthermore, recent studies have revealed profound physical connections among these three diagnostics, establishing a rigorous framework for probing strong-field gravity. In the eikonal limit, a well-established correspondence links QNM frequencies to the properties of the photon sphere and consequently to the shadow radius [43, 44, 45, 46, 47, 48, 49]. Complementing this geometric picture, recent research has demonstrated that the greybody factor plays a pivotal role in governing the excitation amplitude of the ringdown signal, thereby directly linking transmission probabilities to gravitational wave observables [50]. Concurrently, observations suggest that greybody spectra often exhibit greater robustness than quasinormal frequencies under small perturbations of the geometry, offering a complementary diagnostic for the near-horizon structure [51, 52]. More broadly, the scattering interpretation has been deepened to explicitly connect quasinormal frequencies with the transmission and reflection coefficients of the effective potential [53, 54]. This multifaceted correspondence between geometry (shadow), dynamics (QNMs), and scattering (GBFs) now serves as a concrete paradigm for synergistically analyzing black hole physics across different observational channels [55, 56, 57, 58, 59].
@@
Further enriching this landscape, the cloud of strings (CoS) model provides a minimalist geometric framework for describing how one-dimensional extended objects can modify black hole spacetimes. Originally developed by Letelier [89], this model generalizes the concept of a dust cloud to a collection of one-dimensional strings. Physically, it can be interpreted as a macroscopic effective description of topological defects—analogous to cosmic strings—formed during symmetry-breaking phase transitions in the early universe. While not intended to model standard astrophysical accretion flows (such as gas or dust), the CoS serves as a valuable toy model for investigating the gravitational imprints of global anisotropy and extended structures. By treating the background as a continuous fluid with string-like stress, this framework allows one to isolate and study how remnant networks of primordial defects might modify the classical observables in strong-field gravity, distinct from point-particle matter distributions. Recently, studies of black holes immersed in a CoS have attracted growing interest [68, 28, 90, 91, 92, 93].
@@
Figure 1: The lapse function $f(r)$ for diverse parameter configurations. The first panel displays an MCDF-CoS reference case with $a=0.2$, $A=1$, $B=10^{-5}$, $\beta=0.8$, $Q=0.5$, alongside specific limits: MCDF ($a=0$), GCDF ($a=0$, $A=0$, $\beta\neq 1$), CDF ($a=0$, $A=0$, $\beta=1$), and Schwarzschild-de Sitter ($a=0$, $Q=0$). Subsequent panels vary individual parameters while maintaining others at reference values.
@@ II Black hole immersed in a modified Chaplygin-like dark fluid and a cloud of strings
When $a=0$, the black hole metric (1) describes a black hole immersed solely in an MCDF. In the limit $Q\rightarrow 0$, the MCDF lapse function reduces to
@@ IV.3 QNM results and discussion
Our systematic analysis of MCDF-CoS QNMs, computed via third-order WKB and Pöschl-Teller methods (Tables 14, Figs. 46). It is well established that the WKB approximation provides high accuracy for the fundamental modes ($n=0$), particularly in the eikonal regime ($l>n$), but its precision deteriorates rapidly for higher overtones ($n\geq 1$[98, 99]. Consequently, we restrict our quantitative analysis to the fundamental mode ($n=0$) and the first overtone ($n=1$), omitting higher harmonics ($n\geq 2$) where the method becomes unreliable. Generally, we observe strong correlations with the effective potential: higher barriers increase $\mathrm{Re}(\omega)$ (oscillation frequency), while broader barriers decrease $|\mathrm{Im}(\omega)|$ (damping rate).
@@
The overtone structure presented in Table 3 and Fig. 6 compares the fundamental mode ($n=0$) with the first overtone ($n=1$). Consistent with generic black hole behavior, the first overtone exhibits a larger $|\mathrm{Im}(\omega)|$ compared to the fundamental mode, indicating a faster decay rate. For fundamental modes, $|\mathrm{Im}(\omega)|$ decreases and asymptotes with increasing $l$, while $\mathrm{Re}(\omega)$ grows approximately linearly.
@@
Throughout our analysis, the third-order WKB and Pöschl-Teller methods show excellent agreement for the fundamental modes ($n=0$), validating our computational approach in this regime. However, the growing relative error observed for $n=1$ (as shown in Table 3) underscores the necessity of employing higher-precision numerical techniques, such as the continued-fraction method, should a detailed investigation of the higher-overtone spectrum be required in future studies.
@@
Table 3: Dependence of QNM frequencies on the angular quantum number $l$ and overtone number $n$ for the reference black hole configuration: $a=0.2$, $A=1$, $B=10^{-5}$, $\beta=0.8$, $Q=0.5$. The analysis is restricted to the fundamental mode ($n=0$) and the first overtone ($n=1$). $\Delta_{R}\%$ and $\Delta_{I}\%$ denote the relative errors for real and imaginary parts, calculated as $|\omega_{\mathrm{WKB}}-\omega_{\mathrm{P-T}}|/|\omega_{\mathrm{WKB}}|\times 100\%$.
@@
Figure 8: Comparison of greybody factors computed via the WKB method and the QNM correspondence. The left and center panels ($l=1,10$) utilize the higher-order relation [Eq. (45)] including overtone corrections, demonstrating significant improvement in accuracy. The right panel ($l=100$) retains the leading-order approximation, which remains asymptotically exact in the eikonal limit.
@@ V.1 Eikonal limit connection
For massless scalar perturbations, the GBF can be expressed in terms of the quasinormal mode frequencies through the higher-order correspondence relation developed in Ref. [53]. Utilizing the third-order WKB expansion, the exponent function $i\mathcal{K}(\omega)$ in the GBF formula (44) is given by:
$i\mathcal{K}(\omega)=-\frac{\omega^{2}-\mathrm{Re}(\omega_{0})^{2}}{4\mathrm{Re}(\omega_{0})\mathrm{Im}(\omega_{0})}+\Delta_{\mathrm{HO}}(\omega,\omega_{0},\omega_{1}).$ (45)
Here, $\omega_{0}$ and $\omega_{1}$ denote the frequencies of the fundamental mode and the first overtone, respectively. The higher-order correction term $\Delta_{\mathrm{HO}}$, which accounts for the anharmonicity of the effective potential, is explicitly defined in Ref. [53] (see Eq. (4.3) therein). This higher-order formulation significantly improves accuracy compared to the leading-order approximation, particularly in regimes where the potential deviates from a simple barrier shape.
Employing the eikonal QNM approximation (41) for fundamental modes yields
@@
A notable observation is that while the higher-order QNM-GBF correspondence introduced in Eq. (45) demonstrates excellent numerical agreement across the parameter space (Fig. 8), the shadow-GBF relation in Eq. (47) exhibits more pronounced discrepancies. This enhanced sensitivity likely stems from the exponential amplification of small deviations between the actual QNM real part and its geometric optics approximation ($\mathrm{Re}(\omega_{0})-\Omega_{\text{ps}}l$) within the GBF expression. Nevertheless, both formulations consistently underscore the fundamental role of photon sphere geometry in governing wave scattering dynamics. Crucially, our analysis highlights a regime-dependent behavior in achieving this agreement: the inclusion of the higher-order correction term $\Delta_{\mathrm{HO}}$ is essential for low angular momentum modes (e.g., $l=1,10$) where the effective potential is anharmonic, whereas the leading-order approximation naturally converges to the exact result in the eikonal limit (e.g., $l=100$) as the potential barrier becomes harmonic.